OpenStudy (anonymous):

How many solutions does the system have? y=-3x+3 y=X2-3x+5

OpenStudy (pawanyadav):

00

OpenStudy (anonymous):

i'm thinking this have no solution

Directrix (directrix):

I agree.

Directrix (directrix):

You can solve the two equations simultaneously. y=-3x+3 y=X2-3x+5 -3x+3 = X^2-3x+5 3 = x^2 + 5 -2 = x^2 so no real solutions for x.

Directrix (directrix):

@Yishai1

OpenStudy (anonymous):

thanks

Directrix (directrix):

You are welcome.